What magnification is produced by a 1.5 mm focal length microscope objective that is placed at 1.513 mm from the object being viewed? What is the overall magnification of this compound optical microscope if a 7x eyepiece (one that produces a magnification of 7) is used?
Answer #1

given

focal length is f = 1.513 mm

objective that is placed at do = 1.5 mm

using equation

1/di = 1/f- 1/do

= 1/1.513 - 1/1.5

= - 5.728 x 10-3

di = - 174.576 mm

mobj = - 174.576 / - 1.5

mobj = 116.384

so so the linear magnification of the objective lens is116 times

next

M = mobj x Meye

= 116 x 7

the total magnification is used of the microscope is M = 814 times